# Piezoelectric actuation for integrated photonics

### v1

@@ 3.1.2 Tuning Speed
In addition to low power, another key advantage of piezoelectric tuning over thermal tuning is the high tuning speed (i.e. wide actuation bandwidth). The speed mainly relies on three factors including RC constant time, photon lifetime, and mechanical transduction time. Due to the small area of the device and small dielectric constant of AlN (), the capacitor is estimated to be 10 pF which guarantees fast electrical response of 10 ps. The photon lifetime is specific for optical resonators and can be estimated by [216], which is 0.1 ns with optical . This is not a limitation for wide-bandwidth photonic devices such as Mach-Zehnder interferometers (MZI). The high acoustic wave velocity in SiO2 (5900 m/s) helps fast (0.5 ns) building up of the stress field around the waveguide. Based on these observations, sub-nanosecond dynamic response can be expected.
The actuation speed can be characterized by studying the time domain dynamic response of the device as shown in Fig. 10. Under the driving of a small signal (V=2 V) 1 MHz square wave while biasing the laser at the slope of the optical resonance, the output light intensity is modulated [see Fig. 10(a)]. The output follows the input square wave and shows sharp switching between the two voltage levels with high signal to noise ratio (SNR). Furthermore, 5 MHz electrical pulses with 20 ns width is applied in Fig. 10(b). To quantitatively characterize how well the actuator responds to the input electrical signal, the normalized cross-correlation between the input and output signals is calculated as shown in Fig. 10(b) inset. The similarity between the auto-correlation (red) of the input signal itself and the cross-correlation (black) demonstrates ultra-fast actuation beyond the nanosecond. The ultimate tuning speed of the actuator is tested by applying 6 GHz square wave (V=7 V) as in Fig. 10(c). The fact that the output optical modulation (orange) shows oscillation at the same repetition rate of 6 GHz suggests that the piezoelectric actuator is capable of high-speed signal transduction.
@@
The frequency response can be measured using the setup in Fig. 11(a), where the scattering parameters of the actuator are characterized by a vector network analyzer (VNA). The electromechanical S11 measures the reflection of the electrical energy from the actuator while driving with port 1 of the VNA. As shown in Fig. 11(b), mechanical resonances are found from the dips where the electrical energy is converted to the acoustic energy into the device. These resonances will manifest themselves in the modulation of light as measured by the optomechanical S21 response in Fig. 11(c). A series of equally spaced mechanical resonances can be found which align with the dips in S11. These resonances are the HBAR modes formed inside the thick Si substrate which works naturally as an acoustic Fabry-Pérot cavity. As in optics, the spacing between them (also know as the free spectral range, FSR) is inversely proportional to the Si thickness. The first harmonic of the HBAR mode locates at 18.5 MHz. Due to the high quality of the single crystalline Si substrate, a mechanical is obtained, and the high signal contrast between the resonances and anti-resonances severely limits the flatness of the frequency response. Despite of being undesired at low frequencies, these HBAR modes extends up to several GHz which opens a new avenue for efficient acousto-optic modulation as will be discussed later.
Not only the HBAR modes at tens of MHz, the bulk modes of the entire chip at lower frequencies (100s kHz) will also be excited by the piezoelectric actuator as shown in Fig. 11(d). Multiple mechanical modes can be supported including the flexural mode, Lamé mode, and face-shear mode [218]. While the vibration of the flexural mode is mainly out of plane, that of the Lamé and face-shear mode is in plane. The profile of one of the typical Lamé modes is shown in the inset of Fig. 11(d). Both the HBAR and chip modes will distort the transduced signal and need to be mitigated carefully in practical applications.
@@ 3.1.3 Strategies to Improve the Tuning Speed
As mentioned in the last section, the high density of mechanical modes of the chip at MHz frequencies limits the actuation bandwidth of the actuator. In Ref. [52], we investigated a combination of strategies to suppress these chip mechanical modes which resulted in flat response up to 10 MHz. At low frequency, the actuator is relatively small compared with the wavelength of the acoustic wave, so it can be treated as a point source that locally emits acoustic waves. As inspired by Ref. [219] where the flexural modes of a nanomechanical membrane are suppressed by the far-field destructive interference between elastic waves, the flexural mode can be mitigated by differentially driving of two adjacent actuators. As shown in the optical image of the insets in Fig. 12, an auxiliary disk actuator is placed adjacent to the actuator that is on the optical microring resonator. By driving them with the same signal but with phase difference, the excitation of mechanical resonances can be effectively suppressed as shown in the green curve of Fig. 12. However, it is mostly effective for modes below 1 MHz where most flexural modes reside (e.g., modes at 225 and 490 kHz). The in plane Lamé and shear modes are still significant such as the 967 KHz resonance.
@@
To further suppress the in-plane acoustic waves, the entire chip is apodized into a polygon with non-parallel edges to deflect the reflection such that the in-plane acoustic waves cannot resonant. The lower left optical image in Fig. 12 shows one of the apodized photonic chip where the top two corners are intentionally cut to reduce parallel faces. This strategy extends the flatness of the response further to beyond 1 MHz as shown by the red curve in Fig. 12. However, since the light has to be coupled in and out through the left and right edges of the chip, they are still parallel. One of the mechanical mode of the apodized chip at 1.69 MHz is shown in Fig. 12. Finally, by attaching the apodized chip on a piece of carbon tape and then differentially driving the actuators, the flat response is extended up to 10 MHz with variation smaller than 1 dB (blue curve in Fig. 12). This flat response will improve the linearity of piezoelectric actuation for integrated tunable laser, and frequency-modulated continuous-wave (FMCW) LiDAR where any nonlinearity will reduce the SNR as light travels over long distance (100 m) [52]. While these strategies are effective for mechanical modes of the chip, the HBAR modes at tens of MHz that reside vertically in the substrate will require extra solutions. In Ref. [12], we proposed to eliminate the HBAR modes by roughing the backside surface of the Si substrate via isotropic XeF2 etch, and attaching a layer of polyurethane epoxy for absorbing the acoustic waves [220]. However, this method is only effective for HBAR modes at GHz frequencies, where the acoustic wavelength has similar length scale with the backside roughness. On the other hand, since the first harmonic of the HBAR is inversely proportional to the thickness of the substrate, by thinning the Si substrate, the HBAR modes’ frequencies can be increased which can broaden linear actuation bandwidth.
@@ 3.2 Released AlN Actuator
The devices presented in the previous section is solidly mounted on a Si substrate such that the mechanical deformation is limited and the tuning mainly relies on the photoelastic effect. By partially removing the Si beneath the piezoelectric actuator, the actuator gains more degree of freedom to generate mechanical displacement that deforms the shape of the photonic devices. This technique is called “releasing” which has been widely applied in conventional MEMS technologies. The mechanical deformation will modify the photonic devices’ properties in different ways depending on the device design [86]. For example, the gap in an optical directional coupler can be mechanically controlled for tuning the output power ratio in an optical switch [103]. The lengths of the arms in a Mach-Zehnder interferometer can be mechanically stretched for modulating the relative phase [221, 89]. For the optical microring resonators that will be discussed in the following sections, the size of the ring can be expended or shrunk by the piezoelectric actuators which changes the effective cavity length. This pure mechanical effect is usually larger than the photoelastic effect and shows larger tuning, and it is more prominent for smaller devices since it’s proportional to the relative change of the radius .
@@ 3.2.1 DC Tuning
One of the typical released device with AlN piezoelectric actuator is shown in Fig. 14 [222]. A donut shaped actuator is fabricated on the SiO2 cladding which is released by Si isotropic etch under the SiO2 layer. A big center hole together with nine outside holes are opened for the etching gas to react with the Si underneath and form a uniformly released oxide membrane, as can be seen in Fig. 14(b). The free-standing membrane can be seen more clearly from the skewed false-color SEM in Fig. 14(d). The Si3N4 microring is embedded in the SiO2 cladding and located at the inner periphery of the actuator. The waveguide is 6 m away from the inner edge of the oxide cladding [see Fig. 14(c)] which prevents the scattering loss from the sidewall roughness due to the oxide etching. The optical is thus maintained intact after the releasing. The DC tuning of the TE and TM modes is shown in Fig. 14(e), where linear and bi-directional tuning is achieved. The tuning efficiency is around 50 MHz/V for both TE and TM modes respectively, which is over 2 times larger than unreleased device with similar size [12]. Therefore, the tuning can be effectively improved by forming a flexible mechanical structure through releasing.
@@
The fabrication flow for the released device is shown in Fig. 15. The high quality Si3N4 photonic wafer is firstly fabricated with either the subtractive process [223] or the photonic Damascene process [224, 214, 24]. The top surface of the SiO2 cladding undergoes chemical-mechanical polishing (CMP) before the AlN deposition in order to have sub-nanometer surface roughness for better crystalline orientation of the AlN film. Polycrystalline AlN and bottom Mo metal are sputtered on the photonic wafer sequentially in the same vacuum chamber. The AlN is first dry-etched by the reactive-ion-etching (RIE) with gases Cl2 and BCl3 for defining the shape of the AlN actuator. The bottom Mo is dry-etched for patterning the bottom electrode contact pad. The top Al metal is patterned with the standard lift-off process in combination of electron-beam evaporation. In the next step, the releasing holes are made by deep RIE (DRIE) of the SiO2 and Si, which is anisotropic and has good selectivity in the vertical direction as illustrated in Fig. 15(e). Note that the dicing lines between chips are also etched in the same step, which produces smooth chip edge for good lensed fiber to chip edge coupling. In the following, Si isotropic dry etching using SF6 is performed to undercut and suspend the SiO2 membrane. The mechanical polishing in the final step serves to dice the wafer into chips.
@@
For the devices in Fig. 14, the AlN has to be fabricated on flat SiO2 cladding layer for maintaining high quality crystal alignment in the vertical direction of the AlN. This limits the design degree of freedom in the horizontal direction on the SiO2 layer. As shown in Fig. 16, Eichenfield et al. [89] demonstrated a bottom up actuator design where AlN is firstly fabricated before the deposition and etching of the Si3N4 microring resonator. This allows to etch the SiO2 layer cladding to optimize the mechanical deformation. Here, they removed most of the SiO2 in the center and only left a thin wall at the periphery for the optical waveguide [see Fig. 16(b)]. In this way the waveguide has more freedom to move radially for larger relative change of the radius of the microring, as can be seen in Fig. 16(c). On the other hand, the stress is released and is small around the waveguide [see Fig. 16(d)]. Thus, the tuning is mainly from the mechanical deformation of the optical ring.
@@
The tuning at room temperature is experimentally measured as shown in Fig. 16(e). A linear tuning with much higher efficiency of 480 MHz/V is obtained. The efficiency is 2.4 times larger than the unreleased device with similar design which is also studied in Ref. [89]. No hysteresis is observed, illustrating high quality of the AlN film. An ultra-low static power consumption is reported with 8 pW at 2 V, corresponding to power efficiency of 8 fW/MHz. With this low power, they further demonstrated tuning at the cryogenic temperature (7 K) with efficiency of 260 MHz/V. The resistance of the AlN actuator is measured to increase by 40 times at cryogenic temperature which further reduces the leakage current and electrical power dissipation on the order of sub-picowatt. These results prove the suitability of AlN piezoelectric actuator in cryogenic applications which have stringent requirement on power dissipation.
@@ 3.3 Piezoelectric Tuning Using PZT Actuator
Despite of the successful demonstration of piezoelectric tuning using AlN released and unreleased actuator, AlN has a relatively small piezoelectric coefficient which requires high voltage up to 100 V. Although this can be achieved on-chip using charge-discharge pump circuits, they are usually very slow due to their large capacitance. Lead zirconate titanate (PZT), a traditional piezoelectric ceramic that has nearly 20 times larger piezoelectric response, has been widely used in sensors, actuators, and transducers in MEMS community [226]. The successful deposition of PZT thin films and fabrication of PZT actuators on photonic wafers have only been demonstrated in the most recent decade. This section reviews the application of both unreleased and released PZT actuators in the tuning of photonic devices.
@@ 3.3.1 Unreleased PZT tuning of Si Microring Resonator
The PZT tuning of photonic devices is pioneered by U. Levy et al. [106] as shown in Fig. 18. A Si microring resonator with 20 m radius is fabricated on an SOI wafer with 250 nm device layer on a 2-m-thick buried oxide layer (BOX). A 500 nm sol-gel morphotropic PZT (Zr/Ti = 52/48) thin film was spin coated on the photonic wafer, which was then annealed at 650 ∘C for it to crystallize. The PZT is sandwiched between top (100 nm Au) and bottom (100 nm Pt) metal electrodes for applying voltage. Different from previous AlN actuators, the PZT does not require flat top surface to grow crystallization, allowing to pattern the top oxide cladding layer and engineer the distribution of stress. They found that the formation of an oxide cap above the waveguide could concentrate the stress around the oxide corners for enhanced stress-optic tuning, as shown in Fig. 18(c).
Since PZT is also a ferroelectric material, it was electrically poled at 50∘C for 10 min to align the polarization in different domains in the same direction for maximizing its piezoelectric response. The tuning of one optical resonance under different voltages is presented in Fig. 18(d), which shifts to longer wavelength (red tuning) for larger voltage. The resonance is shifted by 34 pm (corresponding to 4.2 GHz change in frequency) at 4 V and by 115 pm (14.4 GHz) at 6 V, which is much larger than previous demonstrations with AlN actuators. Besides the larger piezoelectric response, the larger photoelastic effect of Si also helps to improve the efficiency. Notably, the difference between the spectra at 4 V measured when increasing and decreasing the voltage illustrates the hysteresis behaviour of the PZT actuator. It comes from the ferroelectricity where the polarization can be reoriented by the electric field. Despite of being undesired in applications where linear tuning is required, the hysteresis can be used to realize an optically detected memory system [106]. The tuning speed was tested by applying a 1 kHz square wave to the actuator, and a rise time of 100 s was observed from the optical modulation, which was limited by the photodetector.

### v2

@@ 3.1.2 Tuning Speed
In addition to low power, another key advantage of piezoelectric tuning over thermal tuning is the high tuning speed (i.e. wide actuation bandwidth). The speed mainly relies on three factors including RC constant time, photon lifetime, and mechanical transduction time. Due to the small area of the device and small dielectric constant of AlN (), the capacitor is estimated to be 10 pF which guarantees fast electrical response of 10 ps. The photon lifetime is specific for optical resonators and can be estimated by [227], which is 0.1 ns with optical . This is not a limitation for wide-bandwidth photonic devices such as Mach-Zehnder interferometers (MZI). The high acoustic wave velocity in SiO2 (5900 m/s) guarantees fast (0.5 ns) building up of the stress field around the waveguide, theoretically allowing sub-ps response. Based on these observations, an overall sub-nanosecond dynamic response can be expected.
The actuation speed can be characterized by studying the time domain dynamic response of the device as shown in Fig. 10. Under the driving of a small signal (V=2 V) 1 MHz square wave while biasing the laser at the slope of the optical resonance, the output light intensity is modulated [see Fig. 10(a)]. The output follows the input square wave and shows sharp switching between the two voltage levels with high signal to noise ratio (SNR). Furthermore, 5 MHz electrical pulses with 20 ns width are applied in Fig. 10(b). To quantitatively characterize how well the actuator responds to the input electrical signal, the normalized cross-correlation between the input and output signals is calculated as shown in Fig. 10(b) inset. The similarity between the auto-correlation (red) of the input signal itself and the cross-correlation (black) demonstrates ultra-fast actuation. The ultimate tuning speed of the actuator is tested by applying 6 GHz square wave (V=7 V) as in Fig. 10(c). The fact that the output optical modulation (orange) shows oscillation at the same repetition rate of 6 GHz suggests that the piezoelectric actuator is capable of high-speed signal transduction.
@@
The frequency response can be measured using the setup in Fig. 11(a), where the scattering parameters of the actuator are characterized by a vector network analyzer (VNA). The electromechanical S11 measures the reflection of the electrical energy from the actuator while driving with port 1 of the VNA. As shown in Fig. 11(b), mechanical resonances are found from the dips where the electrical energy is converted to the acoustic energy into the device. These resonances will manifest themselves in the modulation of light as measured by the optomechanical S21 response in Fig. 11(c). A series of equally spaced mechanical resonances can be found which align with the dips in S11. These resonances are the HBAR modes formed inside the thick Si substrate which works naturally as an acoustic Fabry-Pérot cavity. As in optics, the spacing between them (also know as the free spectral range, FSR) is inversely proportional to the Si thickness. The first harmonic of the HBAR mode locates at 18.5 MHz. Due to the high quality of the single crystalline Si substrate, a mechanical is obtained, and the high signal contrast between the resonances and anti-resonances severely limits the flatness of the frequency response and thus the linear transduction bandwidth. The mechanical Q is mainly influenced by the scattering losses occurring at the interfaces of the different layers, which depend on the smoothness of each layer’s surface. Additionally, the damping of the HBAR mode by the bottom chip holder affects the Q, which is related with how the chip is mounted. These loss mechanisms can be exploited to further suppress the mechanical resonances. Despite of being undesired at low frequencies, these HBAR modes extend up to several GHz which opens a new avenue for efficient acousto-optic modulation as will be discussed later.
Bulk mechanical modes of the entire chip at lower frequencies (100s kHz) will also be excited by the piezoelectric actuator as shown in Fig. 11(d). Multiple mechanical modes can be supported including the flexural mode, Lamé mode, and face-shear mode [229]. While the vibration of the flexural mode is mainly out of plane, that of the Lamé and face-shear mode is in plane. The profile of one of the typical Lamé modes is shown in the inset of Fig. 11(d). Both the HBAR and chip modes will distort the transduced signal, in particular in cases where broadband signals such as square and triangle wave modulation is required, and need to be mitigated carefully in practical applications.
@@ 3.1.3 Strategies to Improve the Tuning Speed
As mentioned in the last section, the high density of mechanical modes of the chip at MHz frequencies limits the actuation optomechanical bandwidth. In Ref. [52], we investigated a combination of strategies to suppress these chip mechanical modes which resulted in flat response up to 10 MHz. At low frequency, the actuator is relatively small compared with the wavelength of the acoustic wave, so it can be treated as a point source that locally emits acoustic waves. As inspired by Ref. [230] where the flexural modes of a nanomechanical membrane are suppressed by the far-field destructive interference between elastic waves, the flexural mode can be mitigated by differentially driving of two adjacent actuators. As shown in the optical image of the insets in Fig. 12, an auxiliary disk actuator is placed adjacent to the actuator that is on the optical microring resonator. By driving them with the same signal but with phase difference, the excitation of mechanical resonances can be effectively suppressed as shown in the green curve of Fig. 12. However, it is mostly effective for modes below 1 MHz where most flexural modes reside (e.g., modes at 225 and 490 kHz). The in plane Lamé and shear modes are still significant such as the 967 KHz resonance.
@@
To further suppress the in-plane acoustic waves, the entire chip is apodized into a polygon with non-parallel edges to deflect the reflection such that the in-plane acoustic waves cannot resonate. The lower left optical image in Fig. 12 shows one of the apodized photonic chip where the top two corners are intentionally cut to reduce parallel faces. This strategy extends the flatness of the response further to beyond 1 MHz as shown by the red curve in Fig. 12. However, since the light has to be coupled in and out through the left and right edges of the chip, they are still parallel. One of the mechanical mode of the apodized chip at 1.69 MHz is shown in Fig. 12. Finally, by attaching the apodized chip on a piece of carbon tape and then differentially driving the actuators, the flat response is extended up to 10 MHz with variation smaller than 1 dB (blue curve in Fig. 12). This flat response will improve the linearity of piezoelectric actuation for integrated tunable laser, and frequency-modulated continuous-wave (FMCW) LiDAR where any nonlinearity will reduce the SNR as light travels over long distance (100 m) [52]. While these strategies are effective for mechanical modes of the chip, the HBAR modes at tens of MHz that reside vertically in the substrate will require extra solutions. In Ref. [12], we proposed to eliminate the HBAR modes by roughing the backside surface of the Si substrate via isotropic XeF2 etch, and attaching a layer of polyurethane epoxy for absorbing the acoustic waves [231]. However, this method is only effective for HBAR modes at GHz frequencies, where the acoustic wavelength has similar length scale with the backside roughness. On the other hand, since the first harmonic of the HBAR is inversely proportional to the thickness of the substrate, by thinning the Si substrate, the HBAR modes’ frequencies can be increased which can broaden linear actuation bandwidth.
@@ 3.2 Released AlN Actuator
The devices presented in the previous section is solidly mounted on a Si substrate such that the mechanical deformation is limited and the tuning mainly relies on the photoelastic effect. By partially removing the Si beneath the piezoelectric actuator, the actuator gains more degree of freedom to generate mechanical displacement that deforms the shape of the photonic devices. This technique is called “releasing” which has been widely applied in conventional MEMS technologies. The mechanical deformation will modify the photonic devices’ properties in different ways depending on the device design [88]. For example, the gap in an optical directional coupler can be mechanically controlled for tuning the output power ratio in an optical switch [105]. The lengths of the arms in a Mach-Zehnder interferometer can be mechanically stretched for modulating the relative phase [232, 91]. For the optical microring resonators that will be discussed in the following sections, the size of the ring can be expended or shrunk by the piezoelectric actuators which changes the effective cavity length. This pure mechanical effect is usually stronger than the photoelastic effect and shows larger tuning, and it is more prominent for smaller devices since it’s proportional to the relative change of the radius .
@@ 3.2.1 DC Tuning
One of the typical released device with AlN piezoelectric actuator is shown in Fig. 14 [233]. A donut shaped actuator is fabricated on the SiO2 cladding which is released by Si isotropic etch under the SiO2 layer. A big center hole together with nine outside holes are opened for the etching gas to react with the Si underneath and form a uniformly released oxide membrane, as can be seen in Fig. 14(b). The free-standing membrane can be seen more clearly from the skewed false-color SEM in Fig. 14(d). The Si3N4 microring is embedded in the SiO2 cladding and located at the inner periphery of the actuator. The waveguide is 6 m away from the inner edge of the oxide cladding [see Fig. 14(c)] which prevents the scattering loss from the sidewall roughness due to the oxide etching. The optical is thus maintained intact after releasing. The DC tuning of the TE00 and TM00 modes is shown in Fig. 14(e), where linear and bi-directional tuning is achieved. The tuning efficiency is around 50 MHz/V for both TE00 and TM00 modes respectively, which is over 2 times larger than unreleased device with similar size [12].
@@
The fabrication flow for the released device is shown in Fig. 15. The Si3N4 photonic wafer is firstly fabricated with either the subtractive process [234] or the photonic Damascene process [235, 225, 24]. The top surface of the SiO2 cladding undergoes chemical-mechanical polishing (CMP) before the AlN deposition in order to have sub-nanometer surface roughness for better crystalline orientation of the AlN film. Polycrystalline AlN and bottom Mo metal are sputtered on the photonic wafer sequentially in the same vacuum chamber. The AlN is first dry-etched by the reactive-ion-etching (RIE) with gases Cl2 and BCl3 for defining the shape of the AlN actuator. The bottom Mo is dry-etched for patterning the bottom electrode contact pad. The top Al metal is patterned with the standard lift-off process utilizing electron-beam evaporation. In the next step, the release holes are fabricated by deep RIE (DRIE) of the SiO2 and Si, which is anisotropic and has good selectivity in the vertical direction as illustrated in Fig. 15(e). Note that the dicing lines between chips are also etched in the same step, which produces smooth chip edge for good lensed fiber to chip edge coupling. Subsequently, Si isotropic dry etching using SF6 is performed to undercut and suspend the SiO2 membrane. The mechanical grinding in the final step serves to singularize the wafer into chips.
@@
For the devices in Fig. 14, the AlN has to be fabricated on flat SiO2 cladding layer for maintaining high quality crystal alignment in the vertical direction of the AlN. This limits the design degree of freedom in the horizontal direction on the SiO2 layer. As shown in Fig. 16, Eichenfield et al. [91] demonstrated a buried actuator design where AlN is firstly fabricated before the deposition and etching of the Si3N4 microring resonator. This allows to etch the SiO2 layer cladding to optimize the mechanical deformation. Here, they removed most of the SiO2 in the center and only left a thin wall at the periphery for the optical waveguide [see Fig. 16(b)]. In this way the waveguide has more freedom to move radially for larger relative change of the radius of the microring, as can be seen in Fig. 16(c). On the other hand, the stress is released and is small around the waveguide [see Fig. 16(d)]. Thus, the tuning is mainly from the mechanical deformation of the optical ring.
@@
The tuning at room temperature is experimentally measured as shown in Fig. 16(e). A linear tuning with much higher efficiency of 480 MHz/V is obtained. The efficiency is 2.4 times larger than the unreleased device with similar design which is also studied in Ref. [91]. No hysteresis is observed, illustrating high quality of the AlN film. An ultra-low static power consumption is reported with 8 pW at 2 V, corresponding to power efficiency of 8 fW/MHz. With this low power, they further demonstrated tuning at the cryogenic temperature (7 K) with efficiency of 260 MHz/V. The resistance of the AlN actuator is measured to increase by 40 times at cryogenic temperature which further reduces the leakage current and electrical power dissipation to sub-picowatt levels. These results prove the suitability of AlN piezoelectric actuators in cryogenic applications which have stringent requirement on power dissipation.
@@ 3.3 Piezoelectric Tuning Using PZT Actuator
Despite of the successful demonstration of piezoelectric tuning using AlN released and unreleased actuator, AlN has a relatively small piezoelectric coefficient which requires high voltage up to 100 V. Although this can be achieved on-chip using charge-discharge pump circuits, they are usually very slow due to their large capacitance. PZT, a traditional piezoelectric ceramic that has nearly 20 times larger piezoelectric response, has been widely used in sensors, actuators, and transducers in MEMS community [237]. The successful deposition of PZT thin films and fabrication of PZT actuators on photonic wafers have only been demonstrated in the most recent decade. This section reviews the application of both unreleased and released PZT actuators in the tuning of photonic devices.
@@ 3.3.1 Unreleased PZT tuning of Si Microring Resonator
The PZT tuning of photonic devices is pioneered by U. Levy et al. [108] as shown in Fig. 18. A Si microring resonator with 20 m radius was fabricated on an SOI wafer with 250 nm device layer on a 2-m-thick buried oxide layer (BOX). A 500 nm sol-gel morphotropic PZT (Zr/Ti = 52/48) thin film was spin coated on the photonic wafer, which was then annealed at 650 ∘C for it to crystallize. The PZT is sandwiched between top (100 nm Au) and bottom (100 nm Pt) metal electrodes for applying voltage. Different from previous AlN actuators, the PZT does not require flat top surface to grow polycrystalline, allowing to pattern the top oxide cladding layer and engineer the distribution of stress. They found that the formation of an oxide cap above the waveguide could concentrate the stress around the oxide corners for enhanced stress-optic tuning, as shown in Fig. 18(c).
Since PZT is also a ferroelectric material, it was electrically poled at 50∘C for 10 min to align the polarization in different domains in the same direction for maximizing its piezoelectric response. The tuning of one optical resonance under different voltages is presented in Fig. 18(d), which shifts to longer wavelength (red tuning) for larger voltage. The resonance is shifted by 34 pm (corresponding to 4.2 GHz change in frequency) at 4 V and by 115 pm (14.4 GHz) at 6 V, which is much larger than previous demonstrations with AlN actuators. Besides the larger piezoelectric response, the larger photoelastic effect of Si also helps to improve the tuning efficiency. Notably, the difference between the spectra at 4 V measured when increasing and decreasing the voltage illustrates the hysteresis behaviour of the PZT actuator. It comes from the ferroelectricity where the polarization can be reoriented by the electric field. Despite of being undesired in applications where linear tuning is required, the hysteresis can be used to realize an optically detected memory system [108]. The tuning speed was tested by applying a 1 kHz square wave to the actuator, and a rise time of 100 s was observed from the optical modulation, which was limited by the photodetector.